OpenStudy (anonymous):

When the quadratic equation y = (x − 2)(3x + 1) is written in standard form, what is the value of the coefficient "b"? −5 3 −2 5

OpenStudy (callisto):

y = (x − 2)(3x + 1) = 3x^2 -5x -2 a = 3 b = -5 c= -2